OpenStudy (jollypikachu300):

WILL FAN AND MEDAL HELP Replace ? with =, >, or < to make the statement true. 16 ÷ 4 + 12 ? 4 + 24 ÷ 2 A. < B. > C. =

OpenStudy (jollypikachu300):

@countrygirl1431 @pooja195 help

OpenStudy (anonymous):

\[16\div4=what?\]

OpenStudy (jollypikachu300):

4

OpenStudy (jollypikachu300):

16 divided by 4 equeals 4

OpenStudy (anonymous):

\[4+12=\]

OpenStudy (anonymous):

?

OpenStudy (jollypikachu300):

16

OpenStudy (jollypikachu300):

right

OpenStudy (anonymous):

Good. So what's 4+24?

OpenStudy (anonymous):

?

OpenStudy (jollypikachu300):

29

OpenStudy (jollypikachu300):

@angiecastilloj help

OpenStudy (anonymous):

c

OpenStudy (anonymous):

28 not 29

OpenStudy (anonymous):

both sides equal 16

OpenStudy (anonymous):

28/2=what?

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

>

OpenStudy (anonymous):

No they don't. Check you math. @kelre713

OpenStudy (jollypikachu300):

so c is equal

OpenStudy (anonymous):

its B

OpenStudy (jollypikachu300):

i dont think thats right

OpenStudy (anonymous):

And your not supposed to give out the answer anyway

OpenStudy (jollypikachu300):

ok thx @angiecastilloj

OpenStudy (anonymous):

wwell u seem to be forgetting the biggest rule of math witch is pemdas u divide before u add smart one

OpenStudy (anonymous):

16 ÷ 4 + 12 = 16 4 + 24 ÷ 2 = 14

OpenStudy (jollypikachu300):

oh ok

OpenStudy (anonymous):

24/2=12+4=16 honey now bye

OpenStudy (anonymous):

so its b this is so easy

OpenStudy (anonymous):

fine get it wrong

OpenStudy (jollypikachu300):

@kelre713 shut up